# Data-Efficient Extremum-Seeking Control Using Kernel-Based Function Approximation

### v1

@@ Appendix C Proof of Theorem 15
Step 2: Next, note that a kernel-based update (20) is only performed if $\underline{b}(\theta)>0$ and with $\mu\in\mathbb{R}_{>0}$ chosen such that $\overline{b}(\theta,\mu)\leq 0$ (cf. Algorithm 2). Therefore, it follows from Lemma 13 that
@@
$\Delta V(\theta)\leq-\underline{\rho}_{k,\ell}(V(\theta))$ (71)
at every step, where $\underline{\rho}_{k,\ell}(\cdot):=\min\{\widetilde{\rho}_{k}(\cdot),\dots,\widetilde{\rho}_{k+\ell}(\cdot)\}$ is a function of class $\mathcal{K}$. Similar to Step 1, we conclude from (71) and the comparison lemma that there exists a class-$\mathcal{K}\mathcal{L}$ function $\beta_{2}$ such that
@@
Finally, since $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq V(\theta)$ by Assumption 10 and the definition of $V(\cdot)$ in (40), we obtain from (76) that the optimizer state $\theta$ ultimately converges to the set
$\left\{\theta\in\mathbb{R}^{n_{\theta}}:\lVert\theta\rVert_{\mathcal{C}}\leq\omega_{1}^{-1}\left(\frac{1}{2}\varphi(\gamma+2\delta+\delta_{V})\right)\right\}.$ (77)

### v2

@@ Appendix C Proof of Theorem 15
Step 2: Next, note that a kernel-based update (19) is only performed if $\underaccent{\bar}{b}(\theta)>0$ and with $\mu\in\mathbb{R}_{>0}$ chosen such that $\bar{b}(\theta,\mu)\leq 0$ (cf. Algorithm 1). Therefore, it follows from Lemma 13 that
@@
$\Delta V(\theta)\leq-\underaccent{\bar}{\rho}_{k,\ell}(V(\theta))$ (68)
at every step, where $\underaccent{\bar}{\rho}_{k,\ell}(\cdot):=\min\{\widetilde{\rho}_{k}(\cdot),\dots,\widetilde{\rho}_{k+\ell}(\cdot)\}$ is a function of class $\mathcal{K}$. Similar to Step 1, we conclude from (68) and the comparison lemma that there exists a class-$\mathcal{K}\mathcal{L}$ function $\beta_{2}$ such that
@@
Finally, since $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq V(\theta)$ by Assumption 14, we obtain from (73) that the optimizer state $\theta$ ultimately converges to the set
$\left\{\theta\in\mathbb{R}^{n_{\theta}}\mid\lVert\theta\rVert_{\mathcal{C}}\leq\omega_{1}^{-1}\left(\frac{1}{2}\varphi(\gamma+2\delta+\delta_{V})\right)\right\}.$ (74)